OpenStudy (anonymous):

Solve for x: (x-2)(3x+2)/(x-1)(2x+3) = -1/(x-1) I broke it down so the denominators are the same (x-2)(3x+2)/(x-1)(2x+3) = -(2x+3)/(x-1)(2x+3) Now I am lost, do I distribute?

OpenStudy (anonymous):

(x-2)(3x+2)/(x-1)(2x+3) = -1/(x-1) (x-2)(3x+2) = -1(x-1)(2x+3)/(x-1) (x-2)(3x+2) = -1(2x+3) (x-2)(3x+2) = -2x-3 3x^2 + 2x -6x - 4 = -2x -3 3x^2 -2x - 1 =0 (3x + 1) (x -1)

OpenStudy (anonymous):

@Jlowe did u get it??

OpenStudy (anonymous):

Hot Dog! I thought I was wrong. So it looks like when the instructor said beware of the number you can't use, he was referring to a solution of 1. Thank you